OpenStudy (anonymous):

A recipe that makes 9 servings calls for 4/5 cup of orange juice and 2/3 cup of lemonade. Jeff modifies the recipe so that it can serve 10. How many cups of orange juice does Jeff need? A. 7 1/5 B. 8/9 C. 20/27 D. 4/45

OpenStudy (anonymous):

@omgitsjc @UsukiDoll

OpenStudy (anonymous):

@BSwan

OpenStudy (anonymous):

free medals who ever answers

OpenStudy (anonymous):

@Kainui @

OpenStudy (kainui):

It's not free if I have to answer it. Unless of course you just give me the medal for saying just this.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

is it ok if i follow you

OpenStudy (anonymous):

@omgitsjc

OpenStudy (kainui):

Haha ok, so this is how I think of it. We have amount of orange juice per person currently, but we want the ratio of orange juice to people to remain the same, since each person will get the same amount of orange juice that way. \[\LARGE \frac{4/5 \ cups }{9 \ People}=\frac{OJ}{10 \ People}\] But hey, can't we just solve this for the amount of OJ we need?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

I think the answer is C. but I am not sure.

OpenStudy (anonymous):

@cwrw238 @k

OpenStudy (wolf1728):

No - the answer is B 9 servings requires 4/5 cups of orange juice. 10 servings would require you to increase this by a factor of 10/9 So 4/5 * 10/9 = 40/45 = 8/9 The answer is B